### v1

# Bayesian Stability Selection and Inference on Inclusion Probabilities

Abstract
Stability selection is a versatile framework for structure estimation and variable selection in high-dimensional setting, primarily grounded in frequentist principles. In this paper, we propose an enhanced methodology that integrates Bayesian analysis to refine the inference of inclusion probabilities within the stability selection framework. Traditional approaches rely on selection frequencies for decision-making, often disregarding domain-specific knowledge and failing to account for the inherent uncertainty in the variable selection process. Our methodology uses prior information to derive posterior distributions of inclusion probabilities, thereby improving both inference and decision-making. We present a two-step process for engaging with domain experts, enabling statisticians to elucidate prior distributions informed by expert knowledge while allowing experts to control the weight of their input on the final results. Using posterior distributions, we offer Bayesian credible intervals to quantify uncertainty in the variable selection process. In addition, we highlight how selection frequencies can be uninformative or even misleading when covariates are correlated with each other, and demonstrate how domain expertise can alleviate such issues. Our approach preserves the versatility of stability selection and is suitable for a broad range of structure estimation challenges.
Keywords: Bioinformatics, Bayesian Inference, Model Selection, Stability Selection, Structure Estimation, Variable Selection
1 Introduction
The estimation of discrete structures, including graphs and clusters, and the selection of variables represent long-standing and fundamental problems in statistics (Meinshausen and Bühlmann, 2010). Variable selection is a critical phase in modelling and a crucial step in data pre-processing, which involves modifying and altering raw data to make it suitable for analysis or modelling. Variable selection removes variables that are noisy, redundant, or irrelevant (Tibshirani, 1996; Müller and Welsh, 2010), which is a key challenge in modern data science (Nogueira et al., 2018).
Stability selection, a widely recognised resampling-based variable selection method, determines the selection of variables based on their selection frequencies (Meinshausen and Bühlmann, 2010); which means the frequency with which variables are included in estimated models across multiple resamples.
This is somewhat surprising, given that the discussion section of Meinshausen and Bühlmann (2010) includes some remarks suggesting the potential to explore stability selection from a Bayesian perspective, and despite the substantial attention stability selection has received since its inception, we have found no studies in the literature that integrate it with Bayesian analysis. In this paper, we explore how the Bayesian framework can potentially enhance stability selection by incorporating prior information, offering a comprehensive inferential tool for decision-making.
We demonstrate the methodology using variable selection in the context of linear regression. However, the applications of this approach are not limited to linear regression, as stability selection itself is not limited to this problem. Although beyond the scope of this paper, our framework is applicable for more complex regression-type models.
The idea of using resamples of data to perform variable selection has been used since the past decades. For example, Chen and George (1985) leveraged bootstrap resamples (Efron, 1979) to find important prognostic factors via a stepwise regression procedure in Cox’s proportional hazards model (Cox, 1972). In this context, Rathnayake and Olive (2023) employed bootstrap sampling for variable selection in generalised linear (McCullagh, 2019) and survival models (Cox, 2018).
The stability of a variable selection method denotes the robustness of its preferences for variables across diverse training sets sampled from the same generating distribution (Kalousis et al., 2007) and has been a topic of interest in recent decades (Breiman, 1996b). Evaluation of the stability of a statistical model using random resamples has also been used for decades. For example, Altman and Andersen (1989) applied this approach to Cox’s proportional hazards regression model to identify stable variables, defined as those selected with higher frequencies across multiple resamples.
In the influential work by Meinshausen and Bühlmann (2010), the innovative stability selection framework was introduced, offering a transformative approach for identifying stable variables in structure estimation tasks. The stability selection in Meinshausen and Bühlmann (2010) uses sub-samples with half-size of data to perform variable selection and selects the variables that are consistently identified as important variables across the majority of sub-samples as stable variables. Stability selection can be seen as a meta-algorithm similar to Breiman (1996a) but instead of aggregating the predictions, it aggregates the selections. Meinshausen and Bühlmann (2010) provided an asymptotic upper-bound for the per-family error rate or the expected number of falsely selected variables of the algorithm under the assumptions of exchangeability of the distribution of falsely selected variables and that the selection algorithm is not worse than random guessing.
Shah and Samworth (2013) introduced complementary pairs stability selection, a method where selection frequencies are calculated by counting the number of times a variable is selected in models fitted to complementary 50% sub-samples. Shah and Samworth (2013) introduced more rigorous upper-bounds for the expected number of variables selected with low selection probabilities.
As Holmes noted in the discussion section of Meinshausen and Bühlmann (2010), which is highly relevant to the motivation of this study, statisticians often use crisp variable subset selection to work with data owners to assess the relevance of measured variables and gain a deeper understanding of the dependence structures within the data. Holmes emphasised the importance of providing a formal mechanism for incorporating domain knowledge related to stability selection. Although the results presented by Meinshausen and Bühlmann (2010) depend solely on the data and the algorithm used, the adoption of a Bayesian framework offers a potential way to integrate background knowledge, thus enhancing the results beyond the mere data and algorithmic functions.
The principle of stability selection in Meinshausen and Bühlmann (2010) is based on the repeated sub-sampling of data, using the frequencies with which variables are selected as their selection probabilities to guide decision-making. We advocate for this approach by presenting two main arguments: (1) selection frequencies should be interpreted as sampling probabilities rather than inferential probabilities. As Clayton (2021) explained, “Sampling probabilities go from hypothesis to data: Given an assumption, what will we observe, and how often? Inferential probabilities go from data to hypothesis: Given what we observed, what can we conclude, and with what certainty?”. In this paper, we propose a procedure to enhance sampling probabilities to inferential probabilities, and (2) a sole reliance on frequencies overlooks the importance of background knowledge, thus dismissing the wealth of human understanding in favour of a purely randomised process.
In this paper, we define the inclusion probability as the probability that a variable is incorporated into the estimation of the true underlying model, taking into account both the data and the background knowledge. Our objective is to use selection frequencies to adjust prior probabilities rather than to adopt a position of complete ignorance regarding the variables involved. We further demonstrate, through simulations and analysis of real biological data sets, that correlations between covariates can lead to erratic selection frequencies. In such cases, the incorporation of prior information can improve the decision-making process.
Bayesian variable selection methods, such as Gibbs variable selection (Carlin and Chib, 1995), stochastic search variable selection (George and McCulloch, 1993), adaptive shrinkage (Xu, 2003; Figueiredo, 2003), Bayesian LASSO (Park and Casella, 2008), and reversible jump Markov chain Monte Carlo (Green, 1995), leverage prior information to elucidate parameters, including indicator variables for spike-and-slab models, as well as the location, scale, and rate parameters of coefficients, or the number of variables incorporated into the model (O’Hara and Sillanpää, 2009).
In a Bayesian setting, rarer prior probabilities necessitate greater evidence to surpass the prior and establish an alternative hypothesis. We identified a lack of Bayesian variable selection methods that effectively regulate the extent to which priors dominate the results, thus preventing deterministic outcomes. In this context, we propose a novel and straightforward method for integrating ‘domain expert’ background knowledge into the analysis process, where such expertise is available. This approach improves communication with data owners, allowing them to understand how their expertise is incorporated and the degree to which their input influences the results.
In this paper, we propose the use of the Beta distribution as the prior distribution of inclusion probabilities, which has gained increasing attention in variable selection, as demonstrated by recent studies of Liu and Ročková (2023) and Sparapani et al. (2023). In this regard, we introduce a novel method for determining the parameters of the Beta prior distribution using the experience and knowledge of experts. Specifically, we ask two key questions for each variable under selection, where the first determines the extent to which their background knowledge influences the final results, and the second assesses the perceived importance of the variables. The responses to these questions are then used to elucidate the prior distributions. Although we do not have access to domain experts for our data examples, which limits our ability to directly demonstrate the effectiveness of this approach, we use previously published results to construct informative priors. These priors serve as a proxy for expert knowledge, thereby allowing us to simulate the incorporation of expert insights into our methodology.
To the best of our knowledge, there has been no dedicated research addressing the integration of stability selection and the Bayesian framework, leaving a notable gap in the literature. This study aims to integrate stability selection, which is grounded in frequentist principles, with a novel Bayesian framework to provide inferential inclusion probabilities and Bayesian credible intervals instead of solely relying on sampling frequencies for decision-making. Furthermore, this study poses a question specifically designed to explicitly assess and regulate the degree of subjectivity involved in the final results.
The structure of this paper is organised as follows. Section 2 details the proposed methodology. Section 3 describes the real and synthetic data sets used to apply the methodology. The results of the application of the method to real and synthetic data are presented in Section 4. Finally, Section 5 offers some remarks on Bayesian stability selection, while Section 6 provides a summary and concludes the paper.
2 Methodology
In this section, we introduce a methodology that integrates Bayesian analysis into the stability selection framework. Although we illustrate the proposed method in the context of the stability selection version of Meinshausen and Bühlmann (2010), the proposed methodology is adaptable to other versions of stability selection, for example, as presented in Shah and Samworth (2013) and Beinrucker et al. (2016).
Stability selection, as originally formulated by Meinshausen and Bühlmann (2010), is inherently agnostic to prior information about the variables, relying solely on iterative application of a selection method to different sub-samples of observations. We propose a Bayesian extension to this approach, designed to incorporate expert knowledge into the data-driven selection process. We will outline the specifics of this extension after providing a concise overview of the foundational methodology introduced by Meinshausen and Bühlmann (2010).
2.1 Stability Selection
We consider a data set $\mathcal{D}=\{(\boldsymbol{x}^{\top}_{i},y_{i})\}_{i=1}^{n}$ comprising of a univariate response variable $y_{i}\in\mathbb{R}$ and a $p$-dimensional vector of fixed covariates $\boldsymbol{x}^{\top}_{i}\in\mathbb{R}^{p}$. Typically, when modelling pairs $(\boldsymbol{x}^{\top}_{i},y_{i})$ in the linear regression context, these are assumed to be independent and identically distributed (i.i.d.).
The linear regression model is formally defined as $\boldsymbol{Y}=\beta_{0}+X\boldsymbol{\beta}+\boldsymbol{\varepsilon},$ where $X\in\operatorname{mat}(n\times p)$ denotes the design matrix, $\boldsymbol{Y}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vector representing the univariate response variable, $\beta_{0}\in\mathbb{R}$ denotes the intercept term, $\boldsymbol{\beta}\in\mathbb{R}^{p}$ denotes the $p$-dimensional vector of regression coefficients for the $p$ non-constant covariates, and $\boldsymbol{\varepsilon}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vector of random errors. It is assumed that the components of $\boldsymbol{\varepsilon}$ are i.i.d. and independent of the covariates.
In variable selection, following the terminology of Meinshausen and Bühlmann (2010), it is generally assumed that the covariates can be divided into two distinct groups: the signal group $S=\{k\neq 0\mid\beta_{k}\neq 0\}$ and the noise group $N=\{k\neq 0\mid\beta_{k}=0\}$, where $S\cap N=\emptyset$. A primary goal of a variable selection method is to accurately estimate the signal group $S$.
The stability selection framework requires an appropriate selection method to be implemented within it. One widely used method for variable selection in linear regression setting is the Least Absolute Shrinkage and Selection Operator (LASSO; Tibshirani, 1996), a regularisation technique that employs the $l_{1}$-norm to perform shrinkage on vanishing coefficients. The LASSO estimator is defined as
$\hat{\beta}_{0}(\lambda),\boldsymbol{\hat{\beta}}(\lambda)=\operatorname*{arg\,min}_{\beta_{0}\in\mathbb{R},\boldsymbol{\beta}\in\mathbb{R}^{p}}\left(\|\boldsymbol{Y}-\beta_{0}-X\boldsymbol{\beta}\|_{2}^{2}+\lambda\sum_{k=1}^{p}|\beta_{k}|\right),$ (1)
where $\lambda\in\mathbb{R}^{+}$ is the LASSO regularisation parameter. The set of non-zero coefficients $\hat{S}(\lambda)=\{k\neq 0\mid\hat{\beta}_{k}(\lambda)\neq 0\}$ can be identified using the solution from Equation (1) via convex optimisation.
Meinshausen and Bühlmann (2010) defined the stable set
$\hat{S}^{\text{stable}}=\{j\mid\max_{\lambda\in\Lambda}(\hat{\Pi}_{j}^{\lambda})\geq\pi_{\text{thr}}\};\;j=1,\ldots,p,$
where $\Lambda$ denotes the set of regularisation values, $10<\pi_{\text{thr}}<$ 1 denotes a threshold for decision-making in variable selection, and $\hat{\Pi}_{j}^{\lambda}$ denotes the selection frequency of the $j$th variable given the regularisation parameter $\lambda$.
2.2 Bayesian Stability Selection
We now move forward with the introduction of Bayesian stability selection. In this study, the value $\lambda$ is selected by identifying the largest value of $\lambda$ such that the error remains within one standard error of the minimum error obtained from the cross-validation (Hastie et al., 2009) on all the data $\mathcal{D}$. We use the obtained $\lambda$ value throughout all iterations of the stability selection process. If exploration of different $\lambda$ values is desired, one may employ a grid of $\lambda$ values and repeat the process for each one. Although there are alternative approaches for variable selection in the linear regression model, exploring various methods to determine the regularisation technique and penalty term in stability selection is beyond the scope of this paper.
During each iteration of the stability selection procedure, a random sub-sample comprising half the size of the original data set $\mathcal{D}$ is used. The LASSO model is fitted to this sub-sample, and the binary selection outcomes are recorded as a row in the binary selection matrix $M(\lambda)\in\operatorname{mat}(B\times p)$, where $B$ denotes the total number of sub-samples.
In this context, $M(\lambda)_{bj}=1$ indicates that the $j$th variable is selected as a member of the signal set $\hat{S}(\lambda)$ by applying LASSO to the $b$th sub-sample. Conversely, $M(\lambda)_{bj}=0$ signifies that the $j$th variable is included in the noise set $\hat{N}(\lambda)$ when LASSO is applied to the $b$th sub-sample.
By assuming independence in the selection of variables, similar to Bodinier et al. (2023), we consider a Bernoulli distribution for the selection of each variable in stability selection iterations $M(\lambda)_{bj}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}\text{Bernoulli}(\Pi_{j}^{\lambda});\;j=1,\ldots,p$, where $\Pi_{j}^{\lambda}$ denotes the inclusion probability of the $j$th variable given $\lambda$.
Therefore, the total number of iterations in which the $j$th variable is selected during stability selection, assuming independence of sub-samples, follows a Binomial distribution, that is
$\sum_{b=1}^{B}M(\lambda)_{bj}\sim\text{Binomial}(B,\Pi_{j}^{\lambda});\quad j=1,\ldots,p.$ (2)
To go further, it is necessary to select an appropriate prior distribution for $\Pi_{j}^{\lambda}$. The Beta distribution $\text{Beta}(\alpha,\beta)$ is particularly suitable for modelling probabilities and has support $[0,1]$. Moreover, when the shape parameters of the Beta distribution are not less than $1$, that is, $\alpha,\beta\geq 1$, they are easily interpretable, facilitating the incorporation of background knowledge by directly translating it into a statistical object.
In the Beta distribution, the shape parameters $\alpha$ and $\beta$ offer significant insight into the form and characteristics of the distribution. When $\alpha,\beta>1$, the relative magnitudes of $\alpha$ and $\beta$ determine the skewness of the distribution. Specifically, if $\alpha>\beta$, the corresponding density is left skewed, with the mode shifted toward $1$. In contrast, if $\beta>\alpha$, the density is right skewed, with the mode shifted toward 00. When $\alpha=\beta$, the Beta distribution is symmetric. Importantly, the Beta distribution encompasses the standard Uniform distribution $\mathcal{U}(0,1)$ as a special case when $\alpha=\beta=1$, which represents a non-informative flat prior over the interval $[0,1]$.
Therefore, expert knowledge can be incorporated into the Beta distribution by expressing beliefs through an according choice of the shape parameters $\alpha$ and $\beta$, thereby incorporating prior information into the distribution in a straightforward manner. We will further elaborate on this below.
An additional advantage of the Beta distribution is its conjugacy with the Binomial distribution, which allows for the straightforward estimation of posterior distributions through reparametrisation, thereby obviating the need for complex numerical methods in posterior estimation.
By employing the Beta distribution as prior distribution of inclusion probability for the $j$th variable we have
$\pi(\Pi_{j}^{\lambda})=f(\Pi_{j}^{\lambda};\alpha_{j},\beta_{j})=\frac{({\Pi_{j}^{\lambda}})^{\alpha_{j}-1}(1-\Pi_{j}^{\lambda})^{\beta_{j}-1}}{B(\alpha_{j},\beta_{j})};\quad j=1,\ldots,p\quad\text{and}\quad\alpha_{j},\beta_{j}\geq 1,$ (3)
[…]
Given $n_{j}$ selections of the $j$th variable out of the $B$ iterations, the likelihood function for $\Pi_{j}^{\lambda}$ based on Equation (2) is
$\mathcal{L}(\Pi_{j}^{\lambda}|M(\lambda))=\binom{B}{n_{j}}(\Pi_{j}^{\lambda})^{n_{j}}(1-\Pi_{j}^{\lambda})^{B-n_{j}};\quad j=1,\ldots,p.$ (4)
The posterior distributions can be obtained using the prior distributions in the form of Equation (3) and the information of the rows as observed samples using Equation (4). In other words, each row of $M(\lambda)$ is considered an independent experiment and the prior distribution will be updated in light of the information of the rows. Using Bayes’ theorem, the posterior distribution is proportional to the product of the prior distribution and the likelihood function (Jeffreys, 1934), that is
$\pi(\Pi_{j}^{\lambda}|M(\lambda))\propto(\Pi_{j}^{\lambda})^{\alpha_{j}-1+n_{j}}(1-\Pi_{j}^{\lambda})^{\beta_{j}-1+B-n_{j}},$
which is of the form of the Beta distribution $\Pi_{j}^{\lambda}|M(\lambda)\sim\text{Beta}(\alpha^{\prime}_{j}=\alpha_{j}+n_{j},\beta^{\prime}_{j}=\beta_{j}+B-n_{j})$. Therefore, the posterior density function is as follows
$\pi(\Pi_{j}^{\lambda}|M(\lambda))=f(\Pi_{j}^{\lambda};\alpha_{j},\beta_{j},n_{j},B)=\frac{(\Pi_{j}^{\lambda})^{\alpha_{j}+n_{j}-1}(1-\Pi_{j}^{\lambda})^{\beta_{j}+B-n_{j}-1}}{B(\alpha_{j}+n_{j},\beta_{j}+B-n_{j})}.$ (5)
Equation (5) provides the posterior probability distribution for the inclusion probabilities of variables, allowing for the evaluation of the relevance of variables based on the characteristics of the distribution. The posterior distributions can be used to construct non-asymptotic Bayesian credible intervals for inclusion probabilities, into which the true inclusion probabilities fall into them with the desired probability; providing a clear interpretable tool for communicating the results. In addition, density plots and other distributional characteristics, such as variance, skewness, and kurtosis, can offer a clearer understanding of the selection status of variables.
2.3 Selecting non-informative and informative priors
Here, we propose a method for translating background knowledge into the Beta distributions. As mentioned before, when $\alpha_{j},\beta_{j}\geq 1$, an increase in $\alpha_{j}$ reflects a greater number of selections, thereby shifting the probability distribution towards higher values, whereas an increase in $\beta_{j}$ reflects a greater number of non-selections, shifting the distribution towards lower values. The sum $\gamma_{j}=\alpha_{j}+\beta_{j}$ corresponds to the total number of pseudo-observations considered on the basis of prior information. It is essential that this total does not exceed the number of observed stability selection iterations $B$, to ensure that the prior does not disproportionately influence the inference.
In the absence of prior information, we recommend using $\alpha_{j}=\beta_{j}=1$ as a non-informative prior distribution. As discussed in Section 1, Bayesian variable selection methods integrate background knowledge into the prior distribution by translating expert knowledge into statistical parameters. When such background knowledge is available, we propose a two-step approach to convert this knowledge into the Beta distribution. The first step involves assessing the level of subjectivity included in the final results, while the second step evaluates the perceived importance of the variables under consideration. Specifically, we pose two questions to data owners for each variable subject to selection.
Question 1
Considering that the final results are a synthesis of both your opinions and data-driven insights, what percentage of the final results would you prefer to be influenced by your prior knowledge, with a maximum of $50\%$?
We denote the response to Question 1 by $\tilde{\zeta}_{j}$. For example, $\tilde{\zeta}_{j}=50\%$ indicates that the final result incorporates both $B$ pseudo-observations and $B$ stability selection iterations, with each component contributing $50\%$ of the total weight. The value $\tilde{\zeta}_{j}$ facilitates determining the number of pseudo-observations $\gamma_{j}$ accounted for by the background knowledge. Therefore, the value of $\tilde{\zeta}_{j}$ determines the sum of the prior parameters for the Beta distribution of the $j$th variable as follows
$\frac{\gamma_{j}}{\gamma_{j}+B}=\tilde{\zeta}_{j}.$ (6)
Conceptually, $\lfloor\gamma_{j}\rfloor$ rows are added to $M(\lambda)$ based on the provided prior information.
Question 2
Suppose that you plan to conduct 100 experiments to evaluate the relevance of the $j$th variable. Based on your expertise, what percentage of these experiments do you expect to indicate that the $j$th variable is significant?
We denote the response to Question 2 by $\tilde{\xi}_{j}$, where the value $\tilde{\xi}_{j}$ represents the perceived significance of the $j$th variable. Specifically, $\tilde{\xi}_{j}$ corresponds to the proportion of 1’s in the binary pseudo-observations of the $j$th variable. Based on this, we can calculate the values of $\alpha_{j}$ and $\beta_{j}$ as follows
$\alpha_{j}=\lfloor\tilde{\xi}_{j}\gamma_{j}\rfloor\quad\text{and}\quad\beta_{j}=\lfloor(100\%-\tilde{\xi}_{j})\gamma_{j}\rfloor.$ (7)
Thus, $\tilde{\xi}_{j}$ elucidates the values of the prior parameters, providing insight into the relevance of the $j$th variable based on prior knowledge provided.
Therefore, Equations (6) and (7) offer an interpretable procedure to incorporate background knowledge into statistical distributions. Equation (6) sets the total number of pseudo-observations $\gamma_{j}$, while Equation (7) determines the value of the shape parameters $\alpha_{j}$ and $\beta_{j}$ according to the perceived importance of variables. This approach effectively integrates expert opinions into the Beta distribution treating $B$ as a shared parameter between the two frequentist and Bayesian paradigms to control the degree to which prior knowledge impacts the final results.
To the best of our knowledge, existing Bayesian variable selection methods typically engage in inquiries regarding variable importance, similar to the second question we proposed. This approach seamlessly integrates expert background knowledge into the selection process, thus enhancing the management of the subjectivity-objectivity trade-off.
In Section 4, we present the results obtained from applying the Bayesian stability selection to both synthetic and real data sets, which will be introduced in Section 3.
3 Data Sets
We evaluate Bayesian stability selection using both synthetic data and two real bioinformatics data sets, as detailed below.
3.1 Synthetic Data
We generate data with a sample size of $n=50$ and the number of predictor variables $p=500$, where the predictor variables $\boldsymbol{x}^{\top}_{i}$ are independently drawn from $\mathcal{N}(0,\Sigma)$ where $\Sigma$ is the identity matrix except for the elements $\Sigma_{12}=\Sigma_{34}=\Sigma_{35}=\Sigma_{45}=0.8$ and their symmetric counterparts; therefore, two sets of covariates are correlated: $C_{1}=\{x_{1},x_{2}\}$ and $C_{2}=\{x_{3},x_{4},x_{5}\}$. The response variable is linearly dependent only on the first six predictor variables with coefficient vector
$\boldsymbol{\beta}=(0.9,0.9,0.7,0.7,0.7,1.5,0,\ldots,0)^{\top}$
and the error term $\boldsymbol{\varepsilon}$ is an i.i.d. sample of the zero-centred Normal noise with standard deviation $\sigma=2$. A roughly similar scenario, in which the predictor variables exhibit correlation, is considered in the discussion section of Meinshausen and Bühlmann (2010), by Kirk et al. They showed that in such cases, LASSO tends to select one of the variables from the correlated group at random. As a result, the selection frequencies either favour one variable with a high selection frequency or result in low selection frequencies for all variables of the group. To address this issue, they suggested the use of elastic net (Zou and Hastie, 2005), which combines the LASSO $l_{1}$-penalty with the ridge regression $l_{2}$-penalty (Tikhonov, 1963; Hoerl and Kennard, 1970). We employ the elastic net in this context and demonstrate that the problem in such a case persists. In fact, this is one of the situations in which incorporating background knowledge can assist the algorithm in overcoming data inherent ambiguity.
3.2 Riboflavin Data
The first real data set we use is the well-established ‘riboflavin’ data focused on the production of riboflavin (vitamin B2) from different Bacillus subtilis, provided by Dutch State Mines Nutritional Products, which is available through the hdi R package (Dezeure et al., 2015). The data set comprises a single continuous response variable, representing the logarithm of the riboflavin production rate. In addition, it includes $p=4,088$ covariates, which correspond to the logarithm of the expression levels of $4,088$ bacteria genes. The primary objective of analysing these data is to identify genes that influence riboflavin production, with the ultimate aim of genetically engineering the bacteria to enhance the yield of riboflavin. Data were collected from $n=71$ relatively homogeneous samples, which were repeatedly hybridised during a feed batch fermentation process involving different engineered strains and varying fermentation conditions.
Bühlmann et al. (2014) used stability selection with LASSO and stated that three genes, LYSC_at, YOAB_at, and YXLD_at, have a significant effect on the response variable. Arashi et al. (2021) argued that there are some outlier observations in the riboflavin data and proposed a robust version of the ridge regression, known as rank ridge regression, which they applied to the riboflavin data for gene selection. Their analysis identified $41$ relevant genes, including the three genes previously identified by Bühlmann et al. (2014).
3.3 Affymetrix Rat Genome 230 2.0 Array
The second real data is the ‘Affymetrix Rat Genome 230 2.0 Array’ microarray data set, as introduced by Scheetz et al. (2006). This data set consists of $n=120$ twelve-week-old male rats, with expression levels of nearly 32,000 gene probes recorded for each rat. The objective of this analysis is to identify the probes most strongly associated with the expression level of the TRIM32 probe (1389163_at), which has been linked to the development of Bardet-Biedl syndrome (Chiang et al., 2006), a genetically heterogeneous disorder that affects multiple organ systems, including the retina. In line with the pre-processing steps described by Huang et al. (2008), gene probes with a maximum expression level below the $25$th percentile and those with an expression range smaller than $2$ were excluded. This filtering process resulted in a set of $p=3,083$ gene probes with sufficient expression and variability for further analysis.
Huang et al. (2008) applied adaptive LASSO (Zou, 2006) to analyse this data and identified $19$ gene probes with non-zero effects on the TRIM32 probe. Subsequently, Li et al. (2017) demonstrated that four probes—1389584_at, 1383996_at, 1382452_at, and 1374106_at—are sufficient to model the response variable within a linear framework. With the exception of the third probe, the other probes were also identified by Huang et al. (2008).
4 Results
In this section, we demonstrate the usefulness of Bayesian stability selection using the synthetic and real data introduced in Section 3.
4.1 Synthetic Data
We present the results of applying Bayesian stability selection to synthetic data. We use the elastic net with the mixing parameter $\texttt{alpha}=0.2$. To determine the appropriate $\lambda$ value, we employ a $10$-fold cross-validation using the cv.glmnet function from the glmnet package in R (Friedman et al., 2010) on the entire data set $\mathcal{D}$ and determine the penalty value using the 1-se rule. We generate $100$ different data sets based on the setting described in Section 3 using different random seeds. We set the number of sub-samples $B=100$ and, at the moment, perform the stability selection without using the Bayesian framework.
The average of the selection frequencies of the first six variables that are the relevant variables over the $100$ generated data sets
$\overline{\hat{\Pi}}_{j}=(0.529,0.546,0.604,0.609,0.622,0.540)^{\top}.$
Although the regression coefficient of the variables in $C_{1}$ is higher than those in $C_{2}$, the average selection frequency of the variables in $C_{1}$ is lower than that of $C_{2}$. In addition, although the regression coefficient of $x_{6}$ is double that of the variables in $C_{2}$, its average selection frequency is lower than the variables in $C_{2}$. This suggests that correlated variables, as a group, attract more attention during the selection process compared to uncorrelated variables. In addition, the size of the correlated group directly influences the selection frequencies, with larger groups having higher selection frequencies. Therefore, this simulation reveals that when the relevant variables are highly correlated and the signal-to-noise ratio is small, even the elastic net may fail to address the random selection issue.
Meinshausen and Bühlmann (2010) states that the selection threshold $\pi_{\text{thr}}$ should be selected from the interval $[0.6,0.9]$ therefore, considering $\pi_{\text{thr}}=0.6$, only variables from $C_{2}$ can be considered stable variables. This underscores the importance of incorporating domain knowledge, which can enhance sampling probabilities and help mitigate the ambiguity of the data through prior knowledge and expertise.
Assume that we have selection frequencies, similar to the average selection frequencies mentioned above, that is, $\hat{\Pi}_{j}^{\lambda}=\overline{\hat{\Pi}}_{j}$. Thus, we can say that $n_{j}$ equals 53, 55, 60, 61, 62, and 54 for the first six variables. As discussed in Sections 1 and 2, Bayesian stability selection provides a method for involving data owners’ knowledge in the analysis in a coherent manner. In this example, if we are informed that for these six variables, the domain expert wants half of the final results to be influenced by background knowledge, that is, $\tilde{\zeta_{j}}=50\%$, and these variables are perceived to have an importance of $70\%$, that is, $\tilde{\xi_{j}}=70\%$, using Equations 6 and 7, we derive $\alpha_{j}=70$ and $\beta_{j}=30$. Consequently, applying Equation 5, the mean of the posterior inclusion probabilities for the first six variables $\mathbb{E}[\pi(\Pi_{j}^{\lambda}\mid M(\lambda))]$ is 0.615, 0.625, 0.650, 0.655, 0.660, and 0.620. Therefore, based on the posterior means and adopting $\pi_{\text{thr}}=0.6$, all six relevant variables can be retrieved.
4.2 Riboflavin Data
We apply Bayesian stability selection to riboflavin data. As discussed in Section 3, the data set comprises $p=4,088$ genes and the goal is to identify genes that influence riboflavin production.
Since Arashi et al. (2021) argues that there are some outliers in riboflavin data, we aim to embed a robust selection method within the stability selection framework to evaluate how a robust selector performs on these data, paying attention to outliers. Specifically, we apply Robust Least Angle Regression Selection (RLARS; Khan et al., 2007), a robust linear model selection technique available through the robustHD R package (Alfons, 2021). In the first attempt, non-informative priors are used for all genes and the number of stability selection iterations $B=1,000$. The predictor variables are standardised using the scale function prior to being input into RLARS. Genes with posterior means greater than $0.5$ are presented in Table 1. Based on Table 1, considering the posterior means and using a threshold of $\pi_{\text{thr}}=0.6$, only the gene YCKE_at is deemed stable.
Table 1: List of genes with posterior means exceeding $0.5$
Gene Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
YCKE_at 0.619 0.588 0.649
YOAB_at 0.589 0.558 0.619
YXLD_at 0.525 0.494 0.556
YDAR_at 0.520 0.489 0.551
In the subsequent attempt, we prioritise the $41$ genes detected by Arashi et al. (2021) as effective genes to incorporate background knowledge from their research. We set $\tilde{\zeta}_{j}=50\%$ for these $41$ genes, indicating that this background knowledge have a $50\%$ influence on the final inference of these genes. In addition, we assume $\tilde{\xi}_{j}=70\%$ for all of them, which means that their perceived importance is $70\%$. The hyper-parameters $\tilde{\zeta}_{j}$ and $\tilde{\xi}_{j}$ can be fine-tuned by genetic experts for greater precision. Since $B=1,000$, we have $\alpha_{j}=700$ and $\beta_{j}=300$ for these $41$ genes, while for the remaining genes, we use non-informative priors with $\alpha_{j}=\beta_{j}=1$.
Genes with a posterior mean greater than $0.5$ are presented in Table 2. Based on Table 2, considering the $\pi_{\text{thr}}=0.6$ and posterior mean as the decision benchmark, Bayesian stability selection identified three stable genes using informed priors, which are YCKE_at, YOAB_at, and YXLD_at. The genes YOAB_at and YXLD_at were also identified by Bühlmann et al. (2014); however, the gene LYSC_at is replaced by YCKE_at. We interpret this as evidence that the latter is robust to outlier samples. Furthermore, it can be concluded that not all genes identified by Arashi et al. (2021) are stable when subject to data perturbations.
Table 2: List of genes with posterior means exceeding $0.5$
Gene Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
YCKE_at 0.659 0.639 0.680
YOAB_at 0.644 0.623 0.665
YXLD_at 0.613 0.591 0.634
LYSC_at 0.559 0.538 0.581
YXLE_at 0.544 0.522 0.566
YDAR_at 0.520 0.489 0.551
XHLB_at 0.506 0.484 0.528
The Bayesian credible intervals presented in Tables 1 and 2 have a clear interpretation; given the data and the prior knowledge, the inclusion probabilities of the variables fall into the intervals provided with probability of $0.95$.
4.3 Affymetrix Rat Genome 230 2.0 Array
We apply Bayesian stability selection to the rat genome microarray data introduced in Section 3. Li et al. (2017) identified an outlier in the data set and subsequently removed it, focussing their analysis on the remaining data. However, we employ a robust selection method that accounts for the presence of the outlier. Following the approach used in the riboflavin data, we apply RLARS to identify the key probes across the stability selection iterations. We set the number of stability selection iterations $B=1,000$, and the probes are standardised using the scale function before being incorporated into RLARS. Initially, we use non-informative priors for all probes.
The probes with posterior means greater than 0.3 are presented in Table 3. In particular, none of the four probes listed in Table 3 were identified by Li et al. (2017) or Huang et al. (2008). We interpret this as evidence of the instability of their results when subjected to data perturbation. With a threshold of $\pi_{\text{thr}}=0.6$, no probe can be considered stable. Among the four probes detected by Li et al. (2017), the highest selection frequency corresponds to the probe 1389584_at, with a value of 0.162. Even with $\tilde{\zeta_{j}}=50\%$ and $\tilde{\xi_{j}}=100\%$, this probe has no chance of being classified as stable.
Table 3: List of probes with posterior means exceeding $0.3$
Probe Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
1389457_at 0.528 0.497 0.559
1376747_at 0.523 0.492 0.554
1372928_at 0.399 0.369 0.430
1390539_at 0.345 0.316 0.375
In another approach, we apply the isolation forest model from the isotree R package (Cortes, 2024) to detect outlier samples. The samples GSM130552, GSM130579, GSM130600, GSM130626, GSM130643, and GSM130651 are identified as the 5% most anomalous observations. Consequently, we omitted these samples in this analysis.
The regularisation value $\lambda=0.0371$ is selected by using a 10-fold cross-validation and the 1-se rule on the entire data set $\mathcal{D}$. The probes are standardised using the scale function prior to processing by LASSO. Again, we use non-informative priors.
Probes with posterior means greater than $0.2$ are shown in Table 4. Among these, only the probe 1374106_at was also identified by Li et al. (2017) and Huang et al. (2008). Upon adoption of $\pi_{\text{thr}}=0.6$, none of the probes is considered stable.
Table 4: List of probes with posterior means exceeding $0.2$
Probe Posterior Mean Lower Bound ($95\%$ CI) Upper Bound ($95\%$ CI)
1390539_at 0.360 0.331 0.390
1376747_at 0.218 0.193 0.244
1374106_at 0.210 0.185 0.235
1393727_at 0.207 0.182 0.232
As an example of incorporating prior knowledge, if practitioners have strong beliefs about the importance of probe 1374106_at, based on previous studies such as Huang et al. (2008), these beliefs can be represented as $\tilde{\zeta_{j}}=50\%$ and $\tilde{\xi_{j}}=100\%$. Under these values, the posterior mean of this probe is $0.605$, indicating that it is stable with the adoption of $\pi_{\text{thr}}=0.6$. Clearly, the probe 1390539_at requires less strong beliefs about its importance to be considered stable.
5 Discussion
As Christian Hennig noted in the discussion section of Meinshausen and Bühlmann (2010), it is important in practical applications to incorporate the costs of decisions regarding the inclusion or exclusion of variables into the selection process. A valuable strategy for implementing Bayesian stability selection involves integrating decision costs and conducting risk simulations based on the posterior distributions. By incorporating decision costs, trade-offs can be evaluated between different selection policies and the potential consequences of false positives and false negatives. This approach facilitates a more informed and customised variable selection process, allowing decision makers to simulate various risk scenarios and adjust their selection criteria accordingly.
In case some groups of functionally related variables are of interest, a prior can be imposed on the probability of including at least one member of a group. This approach facilitates the incorporation of background knowledge about a group into the analysis. The corresponding posterior distribution can be derived in a manner analogous to that used for individual variables, offering valuable insight into the selection behaviour of a group as a whole.
Thompson sampling (Thompson, 1933), a well-known Bayesian reinforcement learning algorithm, excels in minimising regret as an agent learns the reward distributions of, for example, a slot machine by sequentially pulling different arms, which is known as the multi-armed bandit problem. Each time the agent selects one or more arms, it updates its beliefs about the reward distributions based on the received rewards, modelled using the Beta distribution. By conceptualising the binary values of the rows of $M(\lambda)$ as the rewards associated with the selection of variables in each iteration, Bayesian stability selection can be seen as training a Thompson-sampling-based agent, where the agent gathers rewards from all arms (that is, variables) at each iteration. Unlike Thompson sampling, where the agent refrains from pulling all arms in every iteration due to associated costs, in the context of Bayesian stability selection, these costs are already accounted for by stability selection, allowing the agent to evaluate all variables simultaneously without such constraints.
One limitation of this paper is the assumption of independence in the selection of variables. This assumption allows us to model the selection status of each variable using an independent Bernoulli distribution. However, a potential improvement would be to account for the interdependencies in variable selection within the modelling process.
6 Conclusion
In this paper, we introduced a Bayesian framework for stability selection that integrates expert knowledge with data-driven insights to provide inference on inclusion probabilities. We specifically argued that selection frequencies should be interpreted as sampling probabilities and proposed a procedure to enhance these into inferential probabilities. The use of Bayesian credible intervals allows for the quantification of uncertainty regarding the selection status of variables. In addition, we demonstrated that selection frequencies can become erratic in the presence of correlated covariates, a common issue when working with biological data under the assumption that statistical correlations reflect molecular interactions.
The conjugacy of the proposed distributions removes the need for numerical estimation of posterior distributions, making Bayesian stability selection an efficient aggregation process following the stability selection iterations. We also proposed a two-step process for communication with experts, enabling the elucidation of prior distributions based on their knowledge while also allowing them to control the degree to which their opinions influence the final results.
Bayesian stability selection remains compatible with existing versions of stability selection, as it is fundamentally built upon the selection matrix, a key component of the original concept introduced by Meinshausen and Bühlmann (2010). As such, it retains the generality of the stability selection framework and can be applied to a wide range of structure estimation tasks, including variable selection.
Competing interests
The authors declare that they have no conflict of interest.
Author contributions statement
Mahdi Nouraie was responsible for drafting the manuscript, the development of the research methodology and for writing the computer code used throughout. Samuel Muller and Connor Smith provided critical feedback on the content of the manuscript, refining the clarity and scope of the manuscript and the computer code.
Data Availability
The riboflavin data set is accessible via the hdi package in R (Dezeure et al., 2015). The rat microarray data can be obtained from the National Center for Biotechnology Information (NCBI) website at www.ncbi.nlm.nih.gov, under accession number GSE5680.
The source code used for the paper is accessible through the following GitHub repository: https://github.com/MahdiNouraie/Bayesian-Stability-Selection.
Acknowledgments
Mahdi Nouraie was supported by the Macquarie University Research Excellence Scholarship (20213605). Samuel Muller was supported by the Australian Research Council Discovery Project Grant (DP230101908).